OpenStudy (anonymous):

4x - 6 = 34 x = __________

OpenStudy (anonymous):

answer is 7

OpenStudy (anonymous):

4x-6=34 4x=34+6 4x=40 x=10

OpenStudy (anonymous):

yea 10

OpenStudy (anonymous):

why 7?

OpenStudy (anonymous):

sorry i added instead of subtract, mistake

OpenStudy (anonymous):

4x- 6= 34 4x=34+6 x=40

OpenStudy (anonymous):

ahh.. ok.. :)

OpenStudy (anonymous):

why 40?

OpenStudy (anonymous):

sorry i mean 10

OpenStudy (anonymous):

ok